# Performance of a SiPM-based, plastic scintillator muon veto prototype for CUPID

### v1

@@ Abstract
CUPID will be a next-generation experiment to search for neutrinoless double beta decay in the inverted mass ordering regime. The reduction of backgrounds in the region of interest is critical to the performance of the experiment. Despite its underground location, muon-induced events will be a non-negligible source of backgrounds for CUPID, and mitigating them is critical to achieving CUPID’s target sensitivity. This necessitates a muon veto for CUPID, so module prototypes were designed, built, and characterized. CUPID’s muon veto system must fit within the physical constraints of the existing infrastructure while maximizing geometrical coverage. A modular system of plastic scintillator panels with embedded plastic wavelength-shifting fibers connected to silicon photomultiplier light detectors is proposed for the CUPID muon tagger. The 10050 cm2 panel prototype presented here exhibited a light yield of () p.e./MeV, with a position reconstruction resolution of 2525 cm2. This design also achieves a detection efficiency of . This paper describes the development, characterization, and testing of prototype modules for the CUPID muon veto. We compare the light yield, uniformity, and position reconstruction potential of different prototype designs.
@@ 1 Muon-induced backgrounds in CUORE and CUPID
Neutrinoless double beta () decay is a theorized ultra-rare decay process that, if observed, would confirm the Majorana nature of neutrinos [1]. Experiments searching for decay, such as the Cryogenic Underground Observatory for Rare Events (CUORE) [2, 3], are currently operating or completed, and the next generation of tonne-scale experiments is in development. CUORE’s successor will be the CUORE Upgrade with Particle Identification (CUPID) [4]. CUORE is an array of 988 TeO2 crystal cryogenic calorimeters located at the Laboratori Nazionali del Gran Sasso (LNGS) [2], planned to continue data taking until 2026. Following CUORE, CUPID will be commissioned in the same experimental infrastructure with upgrades. CUPID will be an array of Li2100MoO4 scintillating crystal cryogenic calorimeters. LNGS is a large underground research facility under the Gran Sasso mountain in central Italy, which provides 3600 meter-water equivalent (m.w.e). of rock shielding. This overburden heavily suppresses the muon flux by about 6 orders of magnitude underground LNGS from approximately at Earth’s surface to [5, 6, 7]. This enables the operation of ultra-low background experiments searching for rare events.
@@
Reducing experimental backgrounds induced by muons is essential for rare event searches like decay. CUPID will fully explore the inverted mass hierarchy region and part of the normal hierarchy region of the effective Majorana mass parameter space. To achieve this, CUPID will employ new technologies and a different isotope (100Mo) with a higher -value ( keV) [8] to decrease the background index (B.I.) in the region of interest (ROI) by about two orders of magnitude compared to CUORE, i.e., from counts/(keV kg y) to counts/(keV kg y) [2, 4]. Although ambitious, this is possible since CUPID will employ a dual-readout technology with heat and light channels to achieve discrimination of alpha events [9], which constitute of the backgrounds in CUPID’s ROI. After alpha rejection, the primary leftover sources of background radiation are radioactive contamination in the detector and its surrounding structures, as well as pile-up events from the two-neutrino double beta (2) decay of 100Mo. However, cosmogenic muons also represent a non-negligible source of backgrounds, even with the attenuation due to the LNGS rock overburden. Muons produce prompt background events by directly depositing energy in the detector, or by producing secondary particles from muon-induced spallation, yielding a rate of approximately 2 events/hour or mHz in CUORE [10]. Muon-induced events yield a background index of counts/(keV kg y) in the ROI for 100Mo decay [4]. Therefore, without any sort of mitigation, muon-induced events alone would saturate all of CUPID’s background budget. An active muon veto will be included in CUPID to reduce this background, with the requirement of rejecting muons in the ROI with an efficiency of at least .
@@
In this work, we will focus on the design and performance of panel prototypes for the muon veto system. The design of the CUPID muon veto system is driven by the need to optimize the muon-tagging efficiency while fitting within the existing spatial constraints of CUORE/CUPID. The background at LNGS for this system is dominated by environmental gamma rays [11], particularly those at MeV from the decay of 208Tl, the gamma ray closest in energy to the muon region of MeV for our design. Another source of background comes from dark noise from the photon-detection devices, which arises from random fluctuations of thermally-generated electrons within a pixel of the photodetectors.
@@
CUORE was built and has been operating without an active muon veto system. To achieve the two orders of magnitude reduction in background required for CUPID, a muon veto system must be added within CUORE’s infrastructure. There are a few potential placements for the panels in the given infrastructure, one of which is shown in figure 1. Performing track reconstruction with the muon veto system is also desirable, but not required. This capability would complement muon track studies and other physics analyses, which rely on events in multiple crystals. Based on these requirements, the key goals of this study were improving light yield, optimizing the design for underground environments, and assessing detection efficiency with different panel configurations.
@@
The proposed muon tagger system consists of single modules of wavelength-shifting (WLS) fibers embedded into a plastic scintillator panel and connected to silicon photomultipliers (SiPMs) as light detectors. The combination of plastic scintillator, WLS fiber, and light detectors is common for muon veto devices [12, 13, 14]. This work describes the development, characterization, and testing of a single module of the muon veto tagger.
@@ 2 Prototype design and fabrication
The panel prototypes described in this paper were designed to study the light yield, background discrimination, and muon tracking capabilities of different single modules. Multiple aspects of the panel’s design and fabrication were tested, such as fiber design, machining techniques, wrapping methods, and coupling to light detectors. Panels embedded with WLS fibers in swirls were chosen to increase the light yield and uniformity by using a high WLS fiber density and minimizing tight turns where light can be lost [15]. In systems using similar materials, increasing the WLS fiber density has been shown to improve light yield [16, 13]. SiPMs were chosen as light detectors to be coupled to the ends of WLS fibers due to space constraints in the CUORE hut for installing muon veto panels. In larger prototypes, multiple spiral patterns were included with the motivation of creating a finer spatial resolution for tracking muons within modules. Modules made from a single panel that self-triggers and a double panel module with a coincidence trigger were also fabricated to test how this could aid in discriminating between muons and gamma backgrounds. Multiple other variables were considered to meet the requirements of these panels and are detailed in this section.
@@
During initial prototyping performed at Yale’s Wright Laboratory, fabrication methods were tested on a 25252.5 cm3 panel. From these tests and the characterization results of this prototype, two different sizes of larger single-module prototypes were developed. The prototypes include a single panel and a double panel module where the panels are of size 100502.5 cm3 and 100501 cm3, respectively (see figure 2). A panel that includes a wavelength-shifting bar and a photomultiplier tube (PMT) was also characterized. This panel was fabricated to be used in a muon veto system at the Low Background Facility at the Lawrence Berkeley National Laboratory (LBNL) and was also studied as a comparison for uniformity and light yield of prototype panels [17]. This panel is about 10 years old, and therefore its performance could also be a proxy for the aging properties of the scintillator material in our design, since they both were procured from the same producer, Eljen Technologies [18].
2.1 Materials selection and fabrication process
All panel prototypes were made with EJ-200 plastic scintillator from Eljen Technologies [18]. 1 mm diameter, round BCF-92 WLS fibers from St. Gobain were embedded in plastic scintillators for Prototypes 1 and 2 [19]. The Small Panel prototype used 1 mm diameter Y-11(200)M WLS fibers from Kuraray [20].
Smaller prototypes were machined and tested at Yale’s Wright Laboratory before scaling to a 10050 cm2 size. To machine grooves for WLS fiber, the plastic scintillator was submerged in a heat bath of soapy water [21]. A programmable CNC machine was used to cut smooth, 4 mm deep grooves. To embed the WLS fiber, UV curable Norland Optical Adhesive 68 was applied where the WLS fiber exited the panel and in small dots along the path of the fiber [22]. Lastly, EJ-500 optical cement was used to permanently attach and optically couple the fiber to the plastic scintillator [23]. During this process, it was found that consistent fiber positioning was difficult to achieve and also critical for light detector alignment with the fibers. A fiber stencil and fiber weight were designed to achieve uniform WLS fiber placement in Prototypes 1 and 2. The stencil and weight secured the fiber at the bottom of the panel grooves so that they could be aligned with the SiPM.
@@ 2.1 Materials selection and fabrication process
To increase the efficiency of light collection, the panels were wrapped in DuPont Tyvek 8740D secured with Tyvek Tape [24, 25]. From tests wrapping small pieces of plastic scintillator, it was found that two layers of DuPont Tyvek 8740D resulted in improvement of light yield over one layer. To block out ambient light, the panels were also wrapped in one layer of aluminum foil and then in a layer of high-performance black masking tape [26]. These layers served as reflectors and light blockers to make the panels light-tight.
@@ 2.2 Prototype designs
The first prototype was made from a single, self-triggered panel of plastic scintillator. The WLS fiber design included 16 quarter turns of a 3 cm bending radius and 4 quarter turns of a 2.5 cm bending radius. This bending radius is smaller than recommended by the manufacturer, however the light loss due to all bends in this pattern is estimated to be below 20% from the specifications [20]. The chosen geometry represents a compromise, which ensures good performance in terms of light yield and uniformity. This layout is shown in figure 2a and is shown under UV light in figure 3.
@@
Two different WLS fiber patterns were tested in Prototypes 1 and 2. This limited the fiber design. The first pattern consisted of 16-quarter turns of a 3 cm bending radius and 4-quarter turns with a 2.5 cm bending radius in a 2525 cm2 panel area. The second included 12-quarter turns of a 5.5 cm bending radius, 4 turns with a 4.5 cm bending radius, and 4 turns with a 2.5 cm radius in a 2525 cm2 panel area. This design is shown in the top view of figure 2c and is shown under UV light in figure 3. The 8 swirls, or "mini-modules," in each prototype read out to a distinct photodetector. We use 13360-3050 Hamamatsu SiPMs [27] for light detection, which have a collection area of 33 mm2, and two fiber ends feed into each SiPM. The bending radius follows the manufacturer’s guidance for the minimum bending radius [28]. The goal of this pattern was to enable a spatial resolution of approximately 2525 cm2 on tagged muons while maintaining a large module size. Larger modules are desirable for installation and operation if a high-light yield can be maintained.
@@
Prototypes 1 and 2 also varied in thickness. The 2.5 cm modules are designed to operate as a single, self-triggered panel, and 1 cm thick modules are stacked together and operated as a double panel module as shown in the bottom view of figure 2c. An event in both the top and bottom panel is needed to trigger the Double Panel Module, which was designed to reduce gamma backgrounds and detect through-going muons with a thinner module. Both self-triggered and coincidence-triggered modules of the same fiber layout were fabricated to test background and muon discrimination in both modes.
LBNL muon panels
@@
In addition to new prototypes, this paper includes results from a muon veto panel used in the Low Background Facility at LBNL. The 76762.5 cm3 LBNL panel was designed to be simple, stable, and easy to operate. It was made from EJ-200 plastic scintillator, which was diamond milled on all sides except a frosted edge that diffuses light. A thin strip of EJ-280 wavelength-shifting plastic was attached to the frosted edge side of the panel. A Hamamatsu R1924A photomultiplier tube was coupled to the end of this WLS plastic bar with an EJ-560 silicon rubber optical interface pad [29]. A schematic of this design is shown in figure 2b. The panel was wrapped in reflecting foil and black vinyl light-tight outer wrap. This panel was manufactured in 2012, and it has been located at the LBNL low background counting facility since fabrication. These panels were shipped from LBNL to Wright Lab and re-characterized in early 2023.
@@
| Prototype | Description/size (cm3) | Trigger Mode | WLS Fiber | Fiber Bending Radii (cm) | Coupling | Readout |
| 1 | 100502.5 | self trigger | BCF-92 | 2.5, 3.0 | mechanical | SiPM |
| 2 | 2 panels, 100501 | coincidence trigger | BCF-92 | 2.5, 4.5, 5.5 | mechanical | SiPM |
| Test Panels | ||||||
| Small Panel | 25252.5 | self trigger | Y-11(200)M | 2.5, 3.0 | air | SiPM |
@@
| LBNL | 76762.5 | self trigger | WLS Plastic Strip (EJ-280) | silicon rubber optical interface (EJ-560) | PMT |
@@ 3 Experimental setup
3.1 SiPM quality assurance testing and coupling
In Prototypes 1 and 2, and the Small Panel, SiPMs were used as light detectors to create a compact, scalable system. These prototypes used 13360-3050 Hamamatsu SiPMs [27]. Low dark noise and crosstalk rates were the primary considerations when choosing the SiPM model. The crosstalk probability of 13360-3050 Hamamatsu SiPMs is at the recommended SiPM operating voltage. At a light yield of 10 photoelectrons (p.e.), the expected dark count rate of the SiPMs is 3 Hz when modeled by the Borel distribution [30].

### v2

@@ Abstract
CUPID will be a next-generation experiment searching for neutrinoless double beta decay in the inverted mass ordering regime. The reduction of backgrounds in the region of interest is critical to the performance of the experiment. Despite its underground location, muon-induced events will be a non-negligible source of background for CUPID, and their mitigation will be critical in reaching CUPID’s target sensitivity. This mitigation will be achieved with a muon veto system, which must fit within the physical constraints of the existing infrastructure while maximizing geometrical coverage. We present the design, construction, and characterization of prototypes for a modular system of plastic scintillator panels with embedded plastic wavelength-shifting fibers connected to silicon photomultipliers for the CUPID muon veto. The 100502.5 cm3 panel prototype presented here exhibited a light yield of () p.e./MeV with a position reconstruction resolution of 2525 cm2. This design also achieves a muon detection efficiency of . We compare the light yield, uniformity, and position reconstruction potential of different prototype designs.
@@ 1 Muon-induced backgrounds in CUORE and CUPID
Neutrinoless double beta decay () is a theorized ultra-rare decay process that, if observed, would confirm the Majorana nature of neutrinos [1]. Experiments searching for , such as the Cryogenic Underground Observatory for Rare Events (CUORE) [2, 3], are currently operating or completed, and the next generation of tonne-scale experiments is in development. CUORE’s successor will be the CUORE Upgrade with Particle Identification (CUPID) [4] experiment. CUORE is an array of 988 TeO2 crystal cryogenic calorimeters located at the Laboratori Nazionali del Gran Sasso (LNGS) [2], planned to continue data taking until 2026. Following CUORE, CUPID will be commissioned in the same experimental infrastructure with upgrades. CUPID will be an array of Li2100MoO4 scintillating crystal cryogenic calorimeters. LNGS is a large underground research facility under the Gran Sasso mountain in central Italy, which provides 3600 meter-water equivalent (m.w.e). of rock shielding. This overburden heavily suppresses the muon flux–by about 6 orders of magnitude–from approximately at Earth’s surface to underground at LNGS [5, 6, 7]. This enables the operation of ultra-low background experiments searching for rare events like .
@@
CUPID will fully explore the inverted mass hierarchy region and part of the normal hierarchy region of the effective Majorana mass parameter space. To achieve this, CUPID will employ new technologies and a different isotope (100Mo) with a higher -value ( keV) [8] to decrease the background index (B.I.) in the region of interest (ROI) by about two orders of magnitude compared to CUORE, i.e., from counts/(keV kg y) to counts/(keV kg y) [2, 4]. To do this, CUPID will employ a dual-readout technology with heat and light channels to achieve discrimination of alpha events [9], which constitute of the backgrounds in CUPID’s ROI. After alpha rejection, the primary leftover sources of background radiation are radioactive contamination in the detector and its surrounding structures, as well as pile-up events from the two-neutrino double beta decay (2) of 100Mo. However, cosmogenic muons also represent a non-negligible source of backgrounds, even with the attenuation from the LNGS rock overburden. Muons produce prompt background events by directly depositing energy in the detector, or by producing secondary particles from muon-induced spallation, yielding a rate of approximately 2 events/hour or mHz in CUORE [10]. Muon-induced events yield a background index of counts/(keV kg y) in the ROI for in 100Mo [4]. This means that without any sort of mitigation, muon-induced events alone would saturate all of CUPID’s background budget. An active muon veto will be included in CUPID to reduce this background, with the requirement that muons in the ROI be rejected with an efficiency of at least .
@@
In this work, we focus on the design and performance of prototype panels for the muon veto system. The design of the CUPID muon veto system is driven by the need to optimize the muon-tagging efficiency while fitting within the existing spatial constraints of the CUORE/CUPID infrastructure. The background at LNGS for this system is dominated by environmental gamma rays [11], particularly those at 2.6 MeV from the decay of 208Tl, the gamma ray closest in energy to the muon signal region of 5 MeV for our design. The prototype panels make use of photodetectors that are themselves a source of background due to dark noise, which arises from random fluctuations of thermally-generated electrons within individual pixels of each photodetector.
@@
CUORE was built and has been operating without an active muon veto system. To achieve the two orders of magnitude reduction in background required for CUPID, a muon veto system must be added within CUORE’s infrastructure. There are a few potential placements for the panels in the given infrastructure, one of which is shown in figure 1. Performing track reconstruction with the muon veto system is also desirable, but not required. This capability would complement muon track studies and other physics analyses that rely on events in multiple crystals. Based on these requirements, the key goals of this study were improving light yield, optimizing the design for underground environments, and assessing detection efficiency with different panel configurations.
@@
The proposed muon veto system consists of single modules of wavelength-shifting (WLS) fibers embedded into a plastic scintillator panel and connected to silicon photomultipliers (SiPMs) as light detectors. The combination of plastic scintillator, WLS fiber, and light detectors is common for muon veto devices [12, 13, 14, 15, 16]. This work describes the development, characterization, and testing of prototype modules for the CUPID muon veto.
@@ 2 Prototype design and fabrication
The prototypes described in this paper were developed to study the light yield, background discrimination, and muon tracking capabilities of different panel designs. Multiple aspects of the panels’ design and fabrication were tested, such as fiber design, machining techniques, wrapping methods, and coupling to light detectors. To increase the light yield and uniformity, panels were embedded with WLS fibers in a swirly pattern that utilizes a high fiber density and minimizes tight turns where light can be lost [17]. In systems using similar materials, increasing the WLS fiber density has been shown to improve light yield [18, 13, 19]. SiPMs were chosen as light detectors to be coupled to the ends of the WLS fibers due to space constraints in the CUORE infrastructure. In larger prototypes, multiple spiral patterns were included with the motivation of creating a finer spatial resolution for tracking muons within modules. A single panel module operated with a self-trigger and a double panel module operated with a coincidence trigger were also fabricated to test how this could aid in discriminating between muons and gamma backgrounds. Multiple other variables were considered to meet the requirements of these panels and are detailed in this section.
@@
During initial prototyping performed at Yale’s Wright Laboratory, fabrication methods were tested on a 25252.5 cm3 panel (the "Small Panel" prototype). From the testing and characterization of this prototype, two different sizes of larger single module prototypes were developed. The prototypes include a single panel module ("Prototype 1") and a double panel module ("Prototype 2") using panels of size 100502.5 cm3 and 100501 cm3, respectively (see figure 2). A panel that includes a wavelength-shifting bar and a photomultiplier tube (PMT) was also characterized. This panel was fabricated to be used in a muon veto system at the Low Background Facility at the Lawrence Berkeley National Laboratory (LBNL) and was also studied to compare uniformity and light yield [20]. This panel is about 10 years old, and therefore its performance could also be a proxy for the aging properties of the scintillator material in our design, since they both were procured from the same producer, Eljen Technologies [21].
2.1 Material selection and fabrication process
@@
All panel prototypes were made with EJ-200 plastic scintillator from Eljen Technologies [21]. 1 mm diameter, round BCF-92 WLS fibers from St. Gobain were embedded in plastic scintillators for Prototypes 1 and 2 [22]. The Small Panel used 1 mm diameter Y-11(200)M WLS fibers from Kuraray [23]. To increase the efficiency of light collection, the panels were wrapped in DuPont Tyvek 8740D secured with Tyvek Tape [24, 25]. From tests wrapping small pieces of plastic scintillator, it was found that two layers of DuPont Tyvek 8740D resulted in improvement of light yield over one layer. To block out ambient light, the panels were also wrapped in one layer of aluminum foil and then in a layer of high-performance black masking tape [26]. These layers serve as reflectors and light blockers to make the panels light-tight.
Before scaling to the full 10050 cm2 surface area, 2525 cm2 panel prototypes were machined and tested at Yale’s Wright Laboratory. To machine grooves for WLS fiber, the plastic scintillator was submerged in a heat bath of soapy water [27]. A programmable Computer Numerical Control (CNC) machine was used to cut smooth, 4 mm deep grooves. To embed the WLS fiber, UV-curable Norland Optical Adhesive 68 was applied where the WLS fiber exited the panel and in small dots along the path of the fiber [28]. Lastly, EJ-500 optical cement was used to permanently attach and optically couple the fiber to the plastic scintillator [29]. During this process, it was found that consistent fiber positioning was difficult to achieve and also critical for the alignment of the SiPM with the fibers. A fiber stencil and counterweight were designed to achieve uniform WLS fiber placement in Prototypes 1 and 2. The stencil and weight secured the fiber at the bottom of the panel grooves so that they could be aligned with the SiPM.
@@ 2.2 Prototype designs
In this section, we describe the designs of the four prototypes of interest to this paper: the Small Panel, Prototype 1, Prototype 2, and the LBNL Panel. A summary of each design is provided in table 1.
@@
The first prototype was made from a single, self-triggered panel of plastic scintillator. The WLS fiber design included 16 quarter turns of a 3 cm bending radius and 4 quarter turns of a 2.5 cm bending radius. The grooves of the design were 4 mm deep, and the straight segments of fiber in the spiral are 2 cm apart. This bending radius is smaller than recommended by the manufacturer, however the light loss due to all bends in this pattern is estimated to be below 20% from the specifications [23]. The chosen geometry was a compromise that ensured good performance in terms of light yield and uniformity. This layout is shown in figure 2 and is shown under UV light in figure 3.
@@
Two different WLS fiber patterns were tested in Prototypes 1 and 2. In Prototype 1 the pattern consists of 16-quarter turns of a 3 cm bending radius and 4-quarter turns with a 2.5 cm bending radius in a 2525 cm2 panel area. This design is shown in the top right of figure 2 and is shown under UV light in figure 3. Prototype 2 uses a different fiber pattern which consists of 12-quarter turns of a 5.5 cm bending radius, 4 turns with a 4.5 cm bending radius, and 4 turns with a 2.5 cm radius in a 2525 cm2 panel area. This design is shown in the bottom right of figure 2. The 8 swirls, or "mini-modules," in each prototype read out to a distinct photodetector. We use 13360-3050PE Hamamatsu SiPMs [30] for light detection, which have a collection area of 33 mm2. Two fiber ends feed into each SiPM. The bending radius follows the manufacturer’s guidance for the minimum bending radius [31]. The goal of this pattern was to enable a spatial resolution of approximately 2525 cm2 on tagged muons while maintaining a large module size. Larger modules are desirable for installation and operation as long as a high light yield can be maintained.
@@
Prototypes 1 and 2 also varied in thickness. The 2.5 cm Prototype 1 module is designed to operate as a single, self-triggered panel, and 1 cm thick Prototype 2 panels are stacked together and operated as a double panel module. An event in both the top and bottom panel is needed to trigger Prototype 2, which was designed to reduce gamma backgrounds and detect through-going muons with a thinner module. Both self-triggered and coincidence-triggered modules of the same fiber layout were fabricated to test background and muon discrimination in both modes.
LBNL Panel
@@
In addition to the new prototypes described above, this paper includes results from a muon veto panel used in the Low Background Facility at LBNL. The 76762.5 cm3 LBNL Panel was designed to be simple, stable, and easy to operate. It was made from EJ-200 plastic scintillator, which was diamond milled on all sides except for a frosted edge that diffuses light. A thin strip of EJ-280 wavelength-shifting plastic was attached to the frosted edge side of the panel. A Hamamatsu R1924A photomultiplier tube was coupled to the end of this WLS plastic bar with an EJ-560 silicon rubber optical interface pad [32]. A schematic of this design is shown in the bottom left of figure 2. The panel was wrapped in reflecting foil and black vinyl light-tight outer wrap. This panel was manufactured in 2012, and it has been located at the LBNL Low Background Counting Facility since fabrication. These panels were shipped from LBNL to Wright Lab and re-characterized in early 2023.
@@
| Prototype | Size (cm3) | Trigger Mode | WLS Fiber | Fiber Bending Radii (cm) | Coupling |
| 1 | 100502.5 | self trigger | BCF-92 | 2.5, 3.0 | mechanical |
| 2 | 2 panels, 100501 | coincidence trigger | BCF-92 | 2.5, 4.5, 5.5 | mechanical |
| Test Panels | |||||
| Small Panel | 25252.5 | self trigger | Y-11(200)M | 2.5, 3.0 | air |
@@
| LBNL | 76762.5 | self trigger | WLS Plastic Strip (EJ-280) | N/A | silicon rubber optical interface (EJ-560) |
@@ 3 Experimental setup
3.1 SiPM coupling and quality assurance
In Prototypes 1 and 2, and the Small Panel, SiPMs were used as light detectors to create a compact, scalable system. These prototypes used 13360-3050 Hamamatsu SiPMs [30, 33]. Low dark noise and crosstalk rates were the primary considerations when choosing the SiPM model. The crosstalk probability of 13360-3050 Hamamatsu SiPMs is at the recommended SiPM operating voltage. At a light yield of 10 photoelectrons (p.e.), the expected dark count rate of the SiPMs is 3 Hz when modeled by the Borel distribution [34].